OpenStudy (anonymous):

Find the circumference of a circle with a diameter of 6 cm.

hartnn (hartnn):

\(Circumference \: C =2\pi r = \pi d \) where d is the diameter!

OpenStudy (anonymous):

im still confused

hartnn (hartnn):

C = pi*d d = diameter = 6 C = 6pi Thats it!

OpenStudy (anonymous):

ohhh thank you

hartnn (hartnn):

welcome ^_^